OpenStudy (anonymous):

Add. 14 1/8 and 26 7/18

OpenStudy (turingtest):

14 1/8=113/8 26 7/18=475/18 LCD of both is 36, so 113/8(4/4)+475/18(2/2)=452/36+950/36=1402/36=701/18 =38 17/18